OpenStudy (anonymous):

Does 3x^2-12=3(x-2)(x+2)?

OpenStudy (angela210793):

3x^2-12=3(x-2)(x+2) 3(x-2)(x+2)=3(x^2-4)=3x^2-12 so yes